An austranaut on a strange planet finds that she can jump a maximum horizontal distance of 15m if her initial speed is 3m/s . Whar is the free fall acceleration on the planet?


1
Expert's answer
2022-05-03T14:21:17-0400
"R_{\max}=\frac{v_0^2}{g}"

"g=\frac{v_0^2}{R_{\max}}=\frac{3^2}{15}=0.6\:\rm m/s^2"